OpenStudy (anonymous):

I tried doing this one on my own, but cant. complete the square. x^2+4x+3=0

OpenStudy (anonymous):

what do the instructions say to do?

OpenStudy (anonymous):

solve the equation by completing the square

OpenStudy (anonymous):

Solve the equation by completing the square. (1) x^2-4x+3=0 First, lets add -3 to both sides and we have: x^2-4x=-3 Now why did we do that? If you look at the left side of the equation only, you will see that its actually a quadratic equation that's missing the "C" term. We are going to choose a new "C" for the left side of the equation and add it to both sides thereby creating a new quadratic equation on the left side that's a perfect square. We'll take 1/2 of the "B" term(1/2)(4) or 2 and square it (2^2=4) and add this to both sides of the equation: x^2-4x+4=-3+4 or (x-2)^2=1 take sqrt of both sides x-2=+or-sqrt(1) or x-2=+or-1 now add 2 to both sidess x=2+or-1 x=2+1=3 and x=2-1=1 ck substitute in (1) 3^2-4(3)+3=0 9-12+3=0 -3+3=0 0=0 1^2-4(1)+3=0 1-4+3=0 -3+3=0 0=0 Hope this helps.

OpenStudy (luigi0210):

well, looks like your problem is solved ^_^'

OpenStudy (anonymous):

or you could do this. 4x^2-x-3=0 4(x^2-x/4-3/4)=0 4(x^2-x/4+1/16-1/16-3/4)=0 4((x-1/8)^2-1/64-3/4)=0 4((x-1/8)^2-49/64)=0 4(x-1/8)^2-4(49/64)=0 4(x-1/8)^2-49/16=0 4(x-1/8)^2=49/16 (x-1/8)^2=49/64 x-1/8 = sqrt(49/64) or x-1/8=-sqrt(49/64) x-1/8 = 7/8 or x-1/8 = -7/8 x = 7/8+1/8 or x = -7/8+1/8 x = 8/8 or x = -6/8 x = 1 or x = -3/4 Answers are x=1 or x=−3 /4

OpenStudy (anonymous):

Im starting to understand, but what would the answer be then?

OpenStudy (luigi0210):

Even though this is completing the square you could check your answers by factoring. x^2+4x+3 (x-3)(x-1) x=1, 3

OpenStudy (luigi0210):

oops

OpenStudy (anonymous):

i told you either x=1 or -3/4

OpenStudy (luigi0210):

*(x+1)(x+3) sorry about that

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

can we do another one?

OpenStudy (anonymous):

yea sure :)

OpenStudy (anonymous):

x^2-10x+5=0

OpenStudy (anonymous):

b=-10 a=1 c=-5

OpenStudy (anonymous):

okay...